a motorist travelled 320 km in five hours.Her average speed for the first 160 km was 80 km per hour.What was her average speed for the second 160 km

Answer by checkley79(3341)

D=RT
160=80T
T=160/80
T=2 HOURS FOR THE FIRST LEG.
5-2=3 HOURS FOR THE SECOND LEG.
160=R*3
R=160/3
R=53.33 MPH IS THE RATE FOR THE SECOND LEG.